OpenStudy (anonymous):

solve each system of equations using sub x+2y=2 y=2x-9

OpenStudy (anonymous):

so x+2(2x-9)=2

OpenStudy (anonymous):

x+4x-18=2

OpenStudy (anonymous):

5x-18=2

OpenStudy (anonymous):

x=4?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so 2(5)-9 =1 y =1

OpenStudy (anonymous):

no 2(4)-9=-1 y=-1

OpenStudy (anonymous):

oh my mistake but some how in back of book it say awnser = 1, -1 and i dont get how they got 1

OpenStudy (anonymous):

the book are not always right maybe they made a little mistake

OpenStudy (anonymous):

ok thank you i was like wow im doing something wrong.